OpenStudy (anonymous):

a pool is shaped like a rectangle with a length 4 times its width w. What is an expression for the distance between opposite corners of the pool?

OpenStudy (anonymous):

Can anyone help?

OpenStudy (anonymous):

\[\text{diagonal}=4w \text{Sec}\left[\text{ArcTan}\left[\frac{1}{4}\right]\right] \]

OpenStudy (anonymous):

Thats the answer? :|

OpenStudy (anonymous):

Notice that no dimensions are given, no numbers like the distance around the pool. There is not enough information to solve for x. The diagonal is a function of w. Another expression: \[\text{diagonal}=\sqrt{w^2 + 16 w^2} \]